OpenStudy (solomonzelman):

I just need a little clue..... continuity/limits Graphing differs from my by-hand conclusion. Should be not hard....

OpenStudy (solomonzelman):

\(\Large\color{blue}{ f(x)=\frac{1}{1+e^{1/x}}\LARGE\color{white}{ \rm │ }}\)

OpenStudy (solomonzelman):

Continuity When I am graphing, I am getting that lim x→0 Does Not Exist But, I am showing though, that x → 0\(\scriptsize\color{slate}{^-}\) 1/x → 0 e \(\scriptsize\color{slate}{^{1/x}}\) → 1 1+ e \(\scriptsize\color{slate}{^{1/x}}\) → 2 So it approaches 1/2 from the left.

OpenStudy (broskishelleh):

What do you need help with?

OpenStudy (solomonzelman):

I am showing, hold on

OpenStudy (broskishelleh):



OpenStudy (broskishelleh):

This is how a graph for it looks, by the way

OpenStudy (solomonzelman):

x→0 \(\scriptsize\color{slate}{^{+}}\) (1/x) → ∞ e^(1/x) → ∞ denominator → ∞ f(x) → 0 x→0 \(\scriptsize\color{slate}{^{-}}\) (1/x) → 0 e^(1/x) → 1 denominator → 2 f(x) → 1/2 however desmos gives me https://www.desmos.com/calculator/mqu6gcfm52

OpenStudy (broskishelleh):

Was I right or wrong, my friend?

OpenStudy (solomonzelman):

wrong, it doesn't approach the same value from left and right side as x approaches zero.

OpenStudy (solomonzelman):

I am just asking that should it be approaching 1/2 from the left side, not 1 ?

OpenStudy (solomonzelman):

e^(1/x) can't → 0

OpenStudy (broskishelleh):

I believe you have to make the approaching of the left 1 by the side, as the calculation basically makes it go a bit higher, like one point, but drops it (Sorry for my bad English)

OpenStudy (solomonzelman):

Make it approach ?

OpenStudy (broskishelleh):

Eh, I am not a very good with speaking math, just doing it

OpenStudy (anonymous):

\[x\to 0\implies \frac{1}{x}\to 0\]??

OpenStudy (solomonzelman):

FROM THE LEFT SIDE, YES

OpenStudy (solomonzelman):

I am just asking that x → 0 \(\scriptsize\color{slate}{ ^{-} }\) 1/x → 0 e\(\normalsize\color{slate}{ ^{1/x} }\) → 1 denominator → 2 f(x) → 1/2

OpenStudy (solomonzelman):

The graph indicates that it approaches 1 from the left though, but e^(1/x) can't be approaching zero, can it ?

OpenStudy (solomonzelman):

e^0 = 1 ?

OpenStudy (solomonzelman):

Re-posting the function, \(\LARGE\color{black}{ f(x)=\frac{1}{1+e^{1/x}} }\)

OpenStudy (anonymous):

depends on the direction \[x\to 0+\implies \frac{1}{x}\to +\infty \implies e^{\frac{1}{x}}\to \infty\implies f\to 0\]

OpenStudy (anonymous):

\[x\to 0^-\implies \frac{1}{x}\to -\infty \implies e^{\frac{1}{x}}\to 0\implies f\to 1\]

OpenStudy (solomonzelman):

x → 0 ^- 1/x → 1/-00.1 → -∞ e^1/x → e^(-1000) → e^1/1000 → e^0 → 1.